TITLE

Thomas D. Schneider^{∗}

version = 1.00 of NAME.tex DATE

Example reference citations: [1, 2, 3]. Claude Shannon wrote this equation [4]:

H = −∑_{i} p_{i}log_{2}p_{i} (bits per symbol)  (1)

Example figure reference: Fig. 1. ⇐Fig 1

References

[1] T. D. Schneider, G. D. Stormo, L. Gold, and A. Ehrenfeucht. Information content of binding sites on nucleotide sequences. J. Mol. Biol., 188:415–431, 1986. http://www.lecb.ncifcrf.gov/˜ toms/paper/schneider1986/.

[2] T. D. Schneider. Information and entropy of patterns in genetic switches. In G. J. Erickson and C. R. Smith, editors, Maximum-Entropy and Bayesian Methods in Science and Engineering, pages 147–154, Dordrecht, The Netherlands, 1988. Kluwer Academic Publishers.

[3] T. D. Schneider and G. D. Stormo. Excess information at bacteriophage T7 genomic promoters detected by a random cloning technique. Nucleic Acids Res., 17:659–674, 1989.

[4] C. E. Shannon. A mathematical theory of communication. Bell System Tech. J., 27:379–423, 623–656, 1948. http://cm.bell-labs.com/cm/ms/what/shannonday/paper.html.

∗National Cancer Institute at Frederick, Laboratory of Experimental and Computational Biology, P. O. Box B, Frederick, MD 21702-1201. (301) 846-5581 (-5532 for messages), fax: (301) 846-5598, email: toms@ncifcrf.gov.

http://www.lecb.ncifcrf.gov/˜ toms/

Figure 1: A caption.

Some example text of the figure.

REMOVE